OpenStudy (anonymous):

if y varies directly as 2p+3.and y=14 when p=2,find y when p=5......(direct variation) can somebody help me please.....i really2 need ur help.....please.....

OpenStudy (anonymous):

y = 2 (2p+3) put p=5

OpenStudy (anonymous):

y=14 when p=2 putting the value of p in this expression: 2p+3, we get 2*2+3 = 7 when the whole expression is 7 (i.e. 2p+3) y=14 which is exactly the double of the value of expression 2*expression = y this is the relation now when p=5 (2*5)+3=13, so y=26, double the expression Hope that makes sense

OpenStudy (anonymous):

@coniepo_ito have you got your answer?